# Mirror Descent Policy Optimisation for Robust Constrained Markov Decision Processes

### v1

@@ Abstract
Safety is an essential requirement for reinforcement learning systems. The newly emerging framework of robust constrained Markov decision processes allows learning policies that satisfy long-term constraints while providing guarantees under epistemic uncertainty. This paper presents mirror descent policy optimisation for robust constrained Markov decision processes (RCMDPs), making use of policy gradient techniques to optimise both the policy (as a maximiser) and the transition kernel (as an adversarial minimiser) on the Lagrangian representing a constrained MDP. In the oracle-based RCMDP setting, we obtain an $\mathcal{O}\left(\frac{1}{T}\right)$ convergence rate for the squared distance as a Bregman divergence, and an $\mathcal{O}\left(e^{-T}\right)$ convergence rate for entropy-regularised objectives. In the sample-based RCMDP setting, we obtain an $\tilde{\mathcal{O}}\left(\frac{1}{T^{1/3}}\right)$ convergence rate. Experiments confirm the benefits of mirror descent policy optimisation in constrained and unconstrained optimisation, and significant improvements are observed in robustness tests when compared to baseline policy optimisation algorithms.
@@ Introduction
Safe RL techniques have been primarily proposed within the constrained Markov decision process (CMDP) formalism (Altman,, 1999). In infinite MDPs, which are characterised by large or continuous state and action spaces, traditional solution techniques include linear programming and dynamic programming with Lagrangian relaxation. Following their empirical performance in high-dimensional control tasks (Lillicrap et al.,, 2015; Mnih et al.,, 2016), policy gradient algorithms have become a particularly popular option for safe RL in control applications (Achiam et al.,, 2017; Yang et al.,, 2020; Ding et al.,, 2020; Liu et al.,, 2021; Xu et al.,, 2021; Paternain et al.,, 2023; Ying et al.,, 2023; Wu et al.,, 2024). As summarised in Table 1, constrained policy gradient techniques often come with provable guarantees regarding global optimum convergence and constraint-satisfaction.
@@
The robustness of mismatches in training and test transition dynamics has been the main subject of interest in the robust MDP (RMDP) framework (Iyengar,, 2005; Nilim and Ghaoui,, 2005), which formulates robust policies by training them on the worst-case dynamics model within a plausible set called the uncertainty set. The framework has led to a variety of theoretically sound policy gradient approaches (Ho et al.,, 2021; Zhang et al.,, 2021; Kumar et al.,, 2023; Kuang et al.,, 2022; Wang and Petrik,, 2024; Li et al.,, 2023; Zhou et al.,, 2023; Wang et al.,, 2023). Recently, the approach has also branched out into the robust constrained MDP (RCMDP) framework (Russel et al.,, 2020, 2023; Wang et al.,, 2022; Bossens,, 2024; Zhang et al.,, 2024; Sun et al.,, 2024), which provides the same robustness notion to CMDPs, where they provide worst-case guarantees on the performance, the constraints, or a combination thereof.
Recent work has shown the benefit of mirror descent policy optimisation techniques in theoretical guarantees. In robust MDPs, state-of-the-art results are obtained for mirror descent based policy gradient (Wang et al.,, 2023), showing an $\mathcal{O}(1/T)$ convergence rate in an oracle-based setting. In constrained MDPs, an $\mathcal{O}(\log(T)/T)$ convergence rate is obtained in the oracle-based setting and an $\tilde{\mathcal{O}}(T^{1/3})$ convergence rate in the sample-based setting.
In terms of coping with continuous state and action spaces, recent approaches include double sampling uncertainty sets, integral probability metrics uncertainty sets, and Wasserstein uncertainty sets (Zhou et al.,, 2023; Abdullah et al.,, 2019; Hou et al.,, 2020).
Table 1: Overview of related work.
(a) Related CMDP policy optimisation algorithms
Algorithm Description Results
CPO (Achiam et al.,, 2017) FIM-approximated trust region solved with convex dual program, line search Monotonically non-decreasing value and constraint-satisfaction throughout the algorithm
RCPO (Tessler et al.,, 2019) primal-dual with projection onto convex set (box constraints), three time-scale (actor,critic,multiplier) convergence to feasible local optimum
PCPO (Yang et al.,, 2020) FIM-approximated trust region with projection to feasible set cf. CPO, and convergence to feasible local optimum
NPG-PD (Ding et al.,, 2020) primal-dual, natural policy gradient ascent with projected subgradient descent on constraint-violation global convergence with $\mathcal{O}(1/T^{1/2})$ for value and constraint-satisfaction (softmax); global convergence rate $\mathcal{O}(1/T^{1/2})$ for value and $\mathcal{O}(1/T^{1/4})$ for constraint-satisfaction (general)
CRPO (Xu et al.,, 2021) value update in interior, constraint reduction in exterior global convergence with $\mathcal{O}(1/T^{1/2})$ for value and constraint-satisfaction (softmax)
PMD-PD (Liu et al.,, 2021) primal-dual, mirror descent, two time-scales (policy, multiplier) global convergence rate $\mathcal{O}(\log(T)/T)$ for value and constraint-satisfaction (softmax)
Sample-based PMD-PD (Liu et al.,, 2021) same as above in the sample-based setting global convergence rate $\tilde{\mathcal{O}}(T^{-1/3})$ for value and constraint-satisfaction (softmax)
(b) Related RMDP and RCMDP policy optimisation algorithms.
Algorithm Description Results
Lyapunov-based RCPG (Russel et al.,, 2023) primal-dual, linear programming for inner problem convergence to local optimum Lyapunov stable policy, Lyapunov RCMDPs
RPD (Wang et al.,, 2022) primal-dual, projected descent with robust TD for inner problem local convergence rate $\mathcal{O}(1/T^{1/4})$, $\delta$-contamination RCMDPs
RC-PO (Sun et al.,, 2024) PCPO, projected gradient descent on parametrised dynamics for separate worst-case value and constraint inner problems cf. CPO, $(s,a)$-rectangular RCMDPs
RMCPMD (Wang and Petrik,, 2024) primal-dual mirror descent approach for policy and transition dynamics global convergence rate $\mathcal{O}(e^{-T})$ to robust value (direct parametrisation); global convergence rate $\mathcal{O}(\frac{1}{T})$ to robust value (softmax), $s$- and $(s,a)$-rectangular RMDPs
Robust PMD-PD (ours) PMD-PD with Lagrangian transition mirror ascent global convergence rate $\mathcal{O}(\frac{1}{T})$ to robust Lagrangian (softmax, squared distance as a Bregman divergence) and $\mathcal{O}(e^{-T})$ to entropy-regularised robust Lagrangian (softmax), $s$- and $(s,a)$-rectangular RCMDPs
Robust Sample-based PMD-PD (ours) same as above in the sample-based setting global convergence rate $\tilde{\mathcal{O}}(\frac{1}{T^{1/3}})$ for robust Lagrangian, value, and constraint-satisfaction (softmax, KL divergence as Bregman divergence), $s$- and $(s,a)$-rectangular RCMDPs)
@@
In this paper, we formulate a policy gradient algorithm for a robust constrained MDP setting, with the aim of providing formal guarantees on the convergence rate. We formulate policy gradient techniques to optimise both the policy (as a maximiser) and the transition kernel (as an adversarial minimiser) on the Lagrangian representing a constrained MDP. For a theoretical analysis, we contribute Robust (Sample-based) PMD, which extends the theory of Liu et al., (2021) and Wang et al., (2023) to the RCMDP setting. We primarily focus on the softmax parametrisation in the theoretical analysis. In the oracle-based RCMDP setting, we obtain an $\mathcal{O}(\frac{1}{T})$ convergence rate for the squared distance as a Bregman divergence, and $\mathcal{O}(e^{-T})$ for entropy-regularised objectives. In the sample-based RCMDP setting, we obtain an $\mathcal{O}(\frac{1}{T^{1/3}})$ convergence rate. For a practical implementation, we formulate MDPO-Robust-Lagrangian (MDPO-Robust-Lag for short), which introduces a robust Lagrangian to Mirror Descent Policy Optimisation (MDPO) (Tomar et al.,, 2022). We evaluate MDPO-Robust-Lagrangian empirically by comparing it to robust-constrained variants of PPO-Lagrangian (Ray et al.,, 2019) and RMCPMD (Wang et al.,, 2023), and find significant improvements in the penalised return in worst-case and average-case test performance on dynamics in the uncertainty set.
@@
The setting of the paper is based on robust constrained Markov Decision Processes (RCMDPs). Formally, an RCMDP is given by a tuple $(\mathcal{S},\mathcal{A},\rho,c_{0},c_{1:m},\mathcal{P},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\rho\in\Delta(\mathcal{S})$ is the starting distribution, $c_{0}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is the cost function, $c_{i}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}\,,i\in[m]$ are constraint-cost functions, $\mathcal{P}$ is the uncertainty set, and $\gamma\in[0,1)$ is a discount factor. In RCMDPs, at any time step $l=0,\dots,\infty$, an agent observes a state $s_{l}\in\mathcal{S}$, takes an action $a_{l}\in\mathcal{A}$ based on its policy $\pi\in\Pi=(\Delta(\mathcal{A}))^{\mathcal{S}}$, and receives cost $c_{0}(s_{l},a_{l})$ and constraint-cost signals $c_{j}(s_{l},a_{l})$ for all $j\in[m]=\{1,\dots,m\}$. Following the action, the next state is sampled from a transition dynamics model $p\in\mathcal{P}$ according to $s_{l+1}\sim p(\cdot|s_{l},a_{l})$. The transition dynamics model $p$ is chosen from the uncertainty set $\mathcal{P}$ to solve a minimax problem. In particular, denoting
$V_{\pi,p}^{j}(\rho):=\mathbb{E}_{s_{0}\sim\rho}\left[\sum_{l=0}^{\infty}\gamma^{t}c_{j}(s_{l},a_{l})|\pi,p\right]$ (1)
@@
$V_{\pi,p}(\rho):=\mathbb{E}_{s_{0}\sim\rho}\left[\sum_{l=0}^{\infty}\gamma^{t}c_{0}(s_{l},a_{l})|\pi,p\right]\,,$ (2)
the goal of the agent is to optimise a minimax objective of the form
$\min_{\pi}\left\{\Phi(\pi):=\sup_{p\in\mathcal{P}}V_{\pi,p}(\rho)\quad\text{s.t.}\quad V_{\pi,p}^{j}(\rho)\leq 0,~{}\forall j\in[m]\right\}\,.$ (3)
The objective has an equivalent unconstrained formulation according to the Lagrangian
$\min_{\pi}\left\{\Phi(\pi):=\sup_{p\in\mathcal{P}}\max_{\lambda\geq 0}L(\pi,p,\lambda):=V_{\pi,p}(\rho)+\sum_{j=1}^{m}\lambda_{j}V_{\pi,p}^{j}(\rho)\right\}\,.$ (4)
If the inner optimisation problem is solved as $p_{k}$ at iteration $k$, this in turn corresponds to an equivalent value function over a traditional MDP (Mei et al.,, 2020; Bossens,, 2024) according to the Lagrangian value function,
@@
To denote the cardinality of sets $\mathcal{S}$ and $\mathcal{A}$, we use the convention $S:=|\mathcal{S}|$ and $A:=|\mathcal{A}|$.
@@
We briefly review the Policy Mirror Descent-Primal Dual (PMD-PD) algorithm (Liu et al.,, 2021) for constrained MDPs. PMD-PD uses natural policy gradient with parametrised softmax policies, which is equivalent to the mirror descent.
At iteration $k\in[K]$ and policy update $t\in[t_{k}]$ within the iteration, PMD-PD defines the regularised state value function,
@@
where $p$ is a fixed transition kernel, $\tilde{\mathbf{c}}_{k}(s_{l},a_{l})=c_{0}(s_{l},a_{l})+\sum_{j=1}^{m}\left(\lambda_{k,j}+\eta_{\lambda}V_{\pi_{k},p_{k}}^{j}\right)c_{j}(s_{l},a_{l})$ and $\alpha\geq 0$ is the regularisation coeffient. The regularisation implements a weighted Bregman divergence which is similar to the KL-divergence. The augmented Lagrangian $\tilde{\lambda}_{k,j}=\lambda_{k,j}+\eta_{\lambda}V_{\pi_{k},p_{k}}^{j}$ is used to improve convergence results by providing a more smooth Lagrangian compared to techniques that clip the Lagrangian. At the end of each iteration $\lambda_{k,j}$ is updated according to $\lambda_{k,j}=\max\left\{\lambda_{k,j}+\eta_{\lambda}V_{\pi_{k+1},p_{k+1}}^{j},-\eta_{\lambda}V_{\pi_{k+1},p_{k+1}}^{j}\right\}$.
@@
While the bold indicates the summation over $1+m$ terms (i.e. the costs and constraint-costs) as in Eq. 5, the tilde notations will be used throughout the text to indicate the use of the augmented Lagrangian multipliers $\tilde{\lambda}_{k}$. However, in Section 4.2 and 4.3, where we do not exploit the augmented property, we state the results generically so they can be applied to both traditional and augmented Lagrangians. Note that since all the costs and constraint-costs range in $[-1,1]$,
@@
Consequently, for any set of multipliers $\lambda\in\mathbb{R}^{m}$, the Lagrangian is bounded by $[-F_{\lambda}/(1-\gamma),F_{\lambda}/(1-\gamma)$ based on the constant
@@
$\tilde{\mathbf{Q}}_{\pi_{k}^{t},p}^{\alpha}(s,a)=\mathbf{Q}_{\pi_{k}^{t}}(s,a)+\frac{\alpha}{1-\gamma}B_{d_{\rho}^{\pi_{k}^{t}}}(\pi_{k}^{t},\pi_{k})\,.$ (14)
This way, PMD-PD applies entropy regularisation with respect to the previous policy as opposed to the uniformly randomized policy used in Cen et al., (2022), allowing PMD-PD to converge to the optimal unregularised policy as opposed to the optimal regularised (and sub-optimal unregularised) policy.
@@
$\pi_{\theta}(a|s)=\frac{e^{\theta_{s,a}}}{\sum_{s,a}e^{\theta_{s,a}}}\,.$ (15)

### v2

@@ Abstract
Safety is an essential requirement for reinforcement learning systems. The newly emerging framework of robust constrained Markov decision processes allows learning policies that satisfy long-term constraints while providing guarantees under epistemic uncertainty. This paper presents mirror descent policy optimisation for robust constrained Markov decision processes (RCMDPs), making use of policy gradient techniques to optimise both the policy (as a maximiser) and the transition kernel (as an adversarial minimiser) on the Lagrangian representing a constrained MDP. Our proposed algorithm obtains an $\tilde{\mathcal{O}}\left(1/T^{1/3}\right)$ convergence rate in the sample-based RCMDP setting. In addition to the RCMDP setting, the paper also contributes an algorithm for approximate gradient descent in the space of transition kernels, which is of independent interest for designing adversarial environments. Experiments confirm the benefits of mirror descent policy optimisation in constrained and unconstrained optimisation, and significant improvements are observed in robustness tests when compared to baseline policy optimisation algorithms.
@@ Introduction
Safe RL techniques have been primarily proposed within the constrained Markov decision process (CMDP) formalism (Altman,, 1999). In infinite MDPs, which are characterised by large or continuous state and action spaces, traditional solution techniques include linear programming and dynamic programming with Lagrangian relaxation. Following their empirical performance in high-dimensional control tasks (Lillicrap et al.,, 2015; Mnih et al.,, 2016), policy gradient algorithms have become a particularly popular option for safe RL in control applications (Achiam et al.,, 2017; Yang et al.,, 2020; Ding et al.,, 2020; Liu et al., 2021b, ; Xu et al.,, 2021; Paternain et al.,, 2023; Ying et al.,, 2023; Wu et al.,, 2024).
@@
The robustness to mismatches in training and test transition dynamics has been the main subject of interest in the robust MDP (RMDP) framework (Iyengar,, 2005; Nilim and Ghaoui,, 2005), which formulates robust policies by training them on the worst-case dynamics model within a plausible set called the uncertainty set. The framework has led to a variety of theoretically sound policy gradient approaches (Ho et al.,, 2021; Zhang et al.,, 2021; Kumar et al.,, 2023; Kuang et al.,, 2022; Wang et al.,, 2024; Li et al.,, 2023; Zhou et al.,, 2023; Wang et al.,, 2023). Recently, the approach has also branched out into the robust constrained MDP (RCMDP) framework (Russel et al.,, 2020, 2023; Wang et al.,, 2022; Bossens,, 2024; Zhang et al.,, 2024; Sun et al.,, 2024), which seeks to provide robustness guarantees to CMDPs.
The framework of RCMDPs is newly emerging and hence the theoretical development lags behind compared to CMDPs and RMDPs. At present, a few challenges have hampered the progress of RCMDP policy optimisation. First, since the framework is based on a worst-case cost function and one or more additional worst-case constraint-cost functions, it is challenging to implement in a single simulator since each of the robust cost functions may turn out to be based on a different transition kernel. This has lead to practical objectives that do not meet the original objective (Russel et al.,, 2020) or which only work for a discrete set of transition kernels (Mankowitz et al.,, 2020). Second, the state-of-the-art CMDP techniques for policy optimisation and RMDP techniques for adversarially optimising the transition dynamics have not yet widely made their way into the RCMDP literature, which has lead to a lack of theoretical guarantees. In RMDPs, state-of-the-art results are obtained for mirror descent based policy gradient algorithm (Wang et al.,, 2024), which applies mirror descent in the policy space and mirror ascent in the transition kernel space. In CMDPs, the PMD-PD algorithm is a related mirror descent based policy gradient algorithm which provides $\mathcal{O}(\log(T)/T)$ convergence rate is obtained in the oracle-based setting and an $\tilde{\mathcal{O}}(1/T^{1/3})$ convergence rate in the sample-based setting (Liu et al., 2021b, ). We then pose the question: using mirror descent based policy optimisation, is it intrinsically more difficult to optimise RCMDPs, or is it possible to obtain similar convergence rates?
@@
With the aim of solving these challenges and providing an affirmative answer to the above question, we draw inspiration from the robust Lagrangian objective (Bossens,, 2024) and the success of mirror descent optimisation in RMDPs and CMDPs (Wang et al.,, 2024; Liu et al., 2021b, ). We show that it is indeed possible to obtain convergence rates similar to Sample-based PMD-PD in the sample-based RCMDP setting. For a theoretical analysis, we contribute Robust Sample-based PMD-PD, which extends the theory of Liu et al., 2021b and Wang et al., (2024) to the RCMDP setting, while focusing on the softmax parametrisation within the sample-based setting. Doing so, we achieve the $\tilde{\mathcal{O}}(1/T^{1/3})$ convergence rate as also obtained by (non-robust) Sample-based PMD-PD. For a practical implementation, we formulate MDPO-Robust-Lagrangian (MDPO-Robust-Lag for short), which introduces a robust Lagrangian to Mirror Descent Policy Optimisation (MDPO) (Tomar et al.,, 2022). We evaluate MDPO-Robust-Lagrangian empirically by comparing it to robust-constrained variants of PPO-Lagrangian (Ray et al.,, 2019) and RMCPMD (Wang et al.,, 2024), and find significant improvements in the penalised return in worst-case and average-case test performance on dynamics in the uncertainty set.
Related work
2.1 Robust constrained objectives
Russel et al., (2020) proposed the robust constrained policy gradient (RCPG) algorithm, which solves the inner problem based on either the robust cost or the robust constraint-cost, which ignores at least one part of the robustness problem in RCMDPs. Solving the RCMDP objective directly, the R3C objective (Mankowitz et al.,, 2020) combines the worst-case costs and worst-case constraint-costs by running distinct simulators with different parametrisations, each being run for one step from the current state. This unfortunately cannot be realised within a single simulator and is only applicable for discrete parametrisations. Additionally, it is also computationally expensive, even more so when there are multiple constraints. Alternatively, the robust Lagrangian objective trades off the cost and different constraint-cost according to whether the constraints are active while requiring only a single simulator with differentiable parameters (Bossens,, 2024).
In this paper, we use the robust Lagrangian objective with multiple constraints and rather than performing a single-step transition kernel update, we perform multi-step updates to the transition kernel at each macro-iteration, in a double-loop manner, to ensure that a near-optimal solution to the inner problem has been found, which in turn allows to provide guarantees on the near-optimality of the policy in solving the robust Lagrangian objective. Assuming that there exists at least one policy that is strictly feasible for all the constraints and all dynamics in the uncertainty set, the policy that optimises the robust Lagrangian will satisfy the worst-case constraint-costs and will, due to complementary slackness, optimise the worst-case value.
2.2 Adversarial transition kernels
Techniques for the design of adversarial transition kernels are of general interest for reinforcement learning. They represent different trade-offs between realism, scalability, computational cost, and sample efficiency. In the context of RMDPs and RCMDPs, they are used in the inner problem, which computes the worst-case transition kernel for the current policy. In general, convex uncertainty and rectangular uncertainty sets are required to make the problem tractable, since non-convex sets and non-rectangular set are in general NP Hard (Wiesemann et al.,, 2013).
The scalability of these techniques is a key challenge. Linear programming based techniques (Russel et al.,, 2020) and techniques based on a small subset of simulators (Mankowitz et al.,, 2020) are not scalable. Instead, gradient-based methods can potentially be performed at scale. In RMDPs, gradient descent techniques for computing the inner problem have been proposed. Unfortunately, previous techniques do not have known guarantees (Abdullah et al.,, 2019; Bossens,, 2024) or provide results for non-rectangular uncertainty sets (which are NP Hard) (Li et al.,, 2023). Recent work has proposed transition mirror ascent (TMA), which comes with an $\mathcal{O}(1/T)$ convergence rate in the oracle-based setting under convex rectangular uncertainty sets, when using fixed learning rates, and $\tilde{\mathcal{O}}(e^{-T})$ convergence rates when using geometric schedules for the learning rate (Wang et al.,, 2024). A similar guarantee was earlier provided with a projected subgradient ascent algorithm, under slightly less general conditions, i.e. when the uncertainty set is a rectangular ball (Kumar et al.,, 2023).
In RCMDPs, the Adversarial RCPG algorithm (Bossens,, 2024) provides a neural representation for transition kernels. Unfortunately, the algorithm does not have known theoretical guarantees. Our work takes note of the oracle-based convergence guarantees of the TMA algorithm of Wang et al., (2024), and extends the approach to design an algorithm for approximate TMA with $\mathcal{O}(1/T^{1/2})$ convergence rate in the sample-based setting. The result corresponds to the same oracle-based rate $\mathcal{O}(e^{-T})$ as in TMA (Wang et al.,, 2024) but accounts for the additional sample complexity from the estimation of the Q-values.
Defining the uncertainty set is an important challenge in the context of RMDPs and RCMDPs. The $\ell_{1}$ norm over the probability space is arguably the most widely used uncertainty set in the RMDP literature. Many other sets, such as $\ell_{p}$ sets, total variation distances, Chi-squared sets, and KL-divergences can be directly related to the $\ell_{1}$ norm. In turn, the bounds of the above sets can be derived from the true data based on relatively simple statistics. However, it is of great interest to construct tighter uncertainty sets with $\ell_{p}$ norms (e.g. Behzadian et al., (2019)) since tighter bounds imply a higher performing policy can be obtained in the true environment (as long as the set still includes the true environment). Beyond their relatively simple statistical derivation, $\ell_{p}$ sets are also fairly easy to project back onto the feasible set. Alternatives include models of uncertainty, suitable for indirect parametrisation and compatible with continuous state-action spaces, for instance the double sampling uncertainty sets, integral probability metrics uncertainty sets, and Wasserstein uncertainty sets (Zhou et al.,, 2023; Abdullah et al.,, 2019; Hou et al.,, 2007). Similarly, $\ell_{p}$ sets can also be applied to indirect parametrisations, which is compatible with the TMA algorithm (Wang et al.,, 2024). While in our main sample-based theorem, directly parametrised $\ell_{p}$ uncertainty sets are used to bound the distance from the nominal, the Lagrangian TMA and Approximate TMA methods can be applied with indirect parametrisations, such as the Gaussian Mixture PTK and the Entropy PTK (Wang et al.,, 2024), as also supported by our experiments.
2.3 Policy optimisation algorithms
Different policy opimisation algorithms have been proposed for RCMDPs and the related frameworks of CMDPs and RMDPs. As summarised in Table 1, the algorithms proposed in the literature work in the primal space, dual space, or a combination thereof, and they provide different guarantees under varying parametrisations.
In the CMDP literature on policy optimisation, the best convergence rates are obtained by PMD-PD (Liu et al., 2021b, ) in both the sample-based and oracle-based setting. Unlike other techniques using regularisation (Cen et al.,, 2022), the algorithm provides unbiased convergence results (i.e. they are expressed as the solution to the unregularised Lagrangian). The technique does not account for variability in the transition kernel, and therefore we extend it to the RCMDP setting.
In the RCMDP literature, policy optimisers either come with no policy convergence guarantees (Russel et al.,, 2020; Mankowitz et al.,, 2020; Bossens,, 2024), basic feasibility and improvement results (Sun et al.,, 2024), or sub-optimal convergence points (Russel et al.,, 2023; Wang et al.,, 2022). Our sample-based theory provides results on par with CMDPs, namely an $\tilde{\mathcal{O}}(1/T^{1/3})$ convergence rate to the robust Lagrangian optimum, a result which is on par with the state-of-the-art PMD-PD algorithm (Liu et al., 2021b, ).
Our algorithm is also practical in the sense that it can be implemented with deep RL techniques. In particular, we use Mirror Descent Policy Optimisation (MDPO) (Tomar et al.,, 2022) as the backbone of our algorithm’s implementation. The technique has been previously studied in traditional MDPs as a highly performing alternative to trust region based deep policy optimisers (e.g. PPO, Schulman et al., (2015) and TRPO, Schulman et al., (2017)). While the clipping implemented in PPO also allows a KL-constrained property, it often loses this property since due to averaging effects the policy is updated even if its clipping ratio is out of bound for some state-action pairs. TRPO computes the Fisher information, which can be expensive, and performs a KL constraint with a reversed order (i.e. $D_{\text{KL}}(\pi_{k},\pi)$ instead of $D_{\text{KL}}(\pi,\pi_{k})$ as used in mirror descent). The MDPO backbone not only leads to a novel algorithm for RCMDPs but also one for CMDPs, which turns out to be highly performant as an alternative to PPO-Lagrangian (Ray et al.,, 2019).
Table 1: Overview of theoretical guarantees for policy optimisation algorithms in CMDPs and RCMDPs. Param indicates the parametrisation. Oracle rate indicates the convergence rate in the oracle setting. Sample rate indicates the convergence rate in the sample-based setting. The notation $\tilde{\mathcal{O}}$ hides logarithmic factors from the big O notation.
Algorithm Algorithm type Param Convergence point Oracle rate Sample rate
CPO (Achiam et al.,, 2017) trust region dual program with backtracking general feasible improvement / /
RCPO (Tessler et al.,, 2019) primal-dual with projection general local CMDP optimum / /
PCPO (Yang et al.,, 2020) trust region with projection general local CMDP optimum / /
NPG-PD (Ding et al.,, 2020) primal-dual natural gradient ascent with projected subgradient descent softmax $\mathcal{O}(1/T^{1/2})$ $\tilde{\mathcal{O}}(1/T^{1/4})$
general $\mathcal{O}(1/T^{1/2})$ cost
$\mathcal{O}(1/T^{1/4})$ constraint
$\tilde{\mathcal{O}}(1/T^{1/4})$ cost
$\tilde{\mathcal{O}}(1/T^{1/6})$ constraint
CRPO (Xu et al.,, 2021) primal softmax global CMDP optimum $\mathcal{O}(1/T^{1/2})$ $\tilde{\mathcal{O}}(1/T^{1/4})$
PMD-PD (Liu et al., 2021b, ) primal-dual mirror descent softmax global CMDP optimum $\tilde{\mathcal{O}}(1/T)$ /
Sample-based PMD-PD (Liu et al., 2021b, ) primal-dual mirror descent softmax global CMDP optimum / $\tilde{\mathcal{O}}(1/T^{1/3})$
Lyapunov-based RCPG (Russel et al.,, 2023) primal-dual, linear programming for inner problem general local RCMDP optimum / /
RPD (Wang et al.,, 2022) primal-dual, projected descent with robust TD stationary point $\mathcal{O}(1/T^{1/4})$ /
RC-PO (Sun et al.,, 2024) trust region with projection general feasible improvement / /
Robust Sample-based PMD-PD (ours) see above softmax global RCMDP optimum / $\tilde{\mathcal{O}}(1/T^{1/3})$
@@ 2.3 Policy optimisation algorithms
The setting of the paper is based on robust constrained Markov Decision Processes (RCMDPs). Formally, an RCMDP is given by a tuple $(\mathcal{S},\mathcal{A},\rho,c_{0},c_{1:m},\mathcal{P},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\rho\in\Delta(\mathcal{S})$ is the starting distribution, $c_{0}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is the cost function, $c_{i}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}\,,i\in[m]$ are constraint-cost functions, $\mathcal{P}$ is the uncertainty set, and $\gamma\in[0,1)$ is a discount factor. To denote the cardinality of sets $\mathcal{S}$ and $\mathcal{A}$, we use the convention $S:=|\mathcal{S}|$ and $A:=|\mathcal{A}|$. In RCMDPs, at any time step $l=0,\dots,\infty$, an agent observes a state $s_{l}\in\mathcal{S}$, takes an action $a_{l}\in\mathcal{A}$ based on its policy $\pi\in\Pi=(\Delta(\mathcal{A}))^{\mathcal{S}}$, and receives cost $c_{0}(s_{l},a_{l})$ and constraint-cost signals $c_{j}(s_{l},a_{l})$ for all $j\in[m]=\{1,\dots,m\}$. Following the action, the next state is sampled from a transition dynamics model $p\in\mathcal{P}$ according to $s_{l+1}\sim p(\cdot|s_{l},a_{l})$. The transition dynamics model $p$ is chosen from the uncertainty set $\mathcal{P}$ to solve a minimax problem. In particular, denoting
$V_{\pi,p}^{j}(\rho):=\mathbb{E}_{s_{0}\sim\rho}\left[\sum_{l=0}^{\infty}\gamma^{l}c_{j}(s_{l},a_{l})|\pi,p\right]$ (1)
@@
$V_{\pi,p}(\rho)=\mathbb{E}_{s_{0}\sim\rho}\left[\sum_{l=0}^{\infty}\gamma^{l}c_{0}(s_{l},a_{l})|\pi,p\right]\,$, (2)
the goal of the agent in a traditional RCMDP is to optimise a minimax objective of the form
$\min_{\pi}\sup_{p\in\mathcal{P}}V_{\pi,p}(\rho)\quad\text{s.t.}\quad\sup_{p_{j}\in\mathcal{P}}V_{\pi,p_{j}}^{j}(\rho)\leq 0,~\forall j\in[m]\,.$ (3)
The objective in Eq. 3 does not represent a single plausible transition kernel for the CMDP since the different suprema are usually met by different transition kernels. In practice, a choice is often made to take the supremum of the traditional cost only (Russel et al.,, 2020). Since this ignores the constraint-costs, one may instead optimise a related problem, namely the robust Lagrangian RCMDP objective (Bossens,, 2024),
$\min_{\pi}\left\{\Phi(\pi):=\sup_{p\in\mathcal{P}}\max_{\lambda\geq 0}\left(V_{\pi,p}(\rho)+\sum_{j=1}^{m}\lambda_{j}V_{\pi,p}^{j}(\rho)\right)\right\}\,$, (4)
which only requires a single simulator and will allow optimisation to focus on constraint-costs if they are violated and on the traditional cost otherwise. If the inner optimisation problem is solved as $p_{k}$ at iteration $k$, this in turn corresponds to an equivalent value function over a traditional MDP (Taleghan and Dietterich,, 2018; Bossens,, 2024) according to the Lagrangian value function,
@@
We briefly review the Policy Mirror Descent-Primal Dual (PMD-PD) algorithm (Liu et al., 2021b, ) for constrained MDPs. PMD-PD uses natural policy gradient with parametrised softmax policies, which is equivalent to the mirror descent.
At iteration $k\in\{0,\dots,K-1\}$ and policy update $t\in\{0,\dots,t_{k}\}$ within the iteration, PMD-PD defines the regularised state value function,
@@
where $p$ is a fixed transition kernel, $\tilde{\mathbf{c}}_{k}(s_{l},a_{l})=c_{0}(s_{l},a_{l})+\sum_{j=1}^{m}\left(\lambda_{k,j}+\eta_{\lambda}V_{\pi_{k},p_{k}}^{j}\right)c_{j}(s_{l},a_{l})$ and $\alpha\geq 0$ is the regularisation coeffient. The regularisation implements a weighted Bregman divergence which is similar to the KL-divergence. The augmented Lagrangian multiplier $\tilde{\lambda}_{k,j}=\lambda_{k,j}+\eta_{\lambda}V_{\pi_{k},p_{k}}^{j}$ is used to improve convergence results by providing a more smooth Lagrangian compared to techniques that clip the Lagrangian multiplier. At the end of each iteration $\lambda_{k,j}$ is updated according to $\lambda_{k,j}=\max\left\{\lambda_{k,j}+\eta_{\lambda}V_{\pi_{k+1},p_{k+1}}^{j},-\eta_{\lambda}V_{\pi_{k+1},p_{k+1}}^{j}\right\}$.
@@
While the bold indicates the summation over $1+m$ terms (i.e. the costs and constraint-costs) as in Eq. 5, the tilde notations will be used throughout the text to indicate the use of the augmented Lagrangian multipliers $\tilde{\lambda}_{k}$. Note that since all the costs and constraint-costs range in [1,1][-1,1],
@@
Consequently, for any set of multipliers $\lambda\in\mathbb{R}^{m}$, the Lagrangian is bounded by $F_{\lambda}/(1-\gamma),F_{\lambda}/(1-\gamma)$ based on the constant
@@
$\tilde{\mathbf{Q}}_{\pi_{k}^{t},p}^{\alpha}(s,a)=\tilde{\mathbf{Q}}_{\pi_{k}^{t}}(s,a)+\frac{\alpha}{1-\gamma}B_{d_{\rho}^{\pi_{k}^{t}}}(\pi_{k}^{t},\pi_{k})\,.$ (14)
In this way, PMD-PD applies entropy regularisation with respect to the previous policy as opposed to the uniformly randomized policy used in Cen et al., (2022), allowing PMD-PD to converge to the optimal unregularised policy as opposed to the optimal regularised (and sub-optimal unregularised) policy.
@@
$\pi_{\theta}(a|s)=\frac{e^{\theta_{s,a}}}{\sum_{a^{\prime}}e^{\theta_{s,a^{\prime}}}}\,.$ (15)